What is the probability of observing a a value more then .75 from a standard normal distribution?
I got 1-.7734
What is the probability of observing a number less then 5 or more than 10 from a normal distribution with mean 9 and standard deviation 2?

Answer by stanbon(75887)

What is the probability of observing a a value more then .75 from a standard normal distribution?
I got 1-.7734 = 0.2266
-----------------------------------
What is the probability of observing a number less then 5 or more than 10 from a normal distribution with mean 9 and standard deviation 2?
---
z(5) = (5-9)/2 = -2
z(10) = (10-9)/2 = 1/2
-------
P(x < -2) or P(x > 1/2)
= 0.0228 + 0.3985
---
= 0.3303
==================
Cheers,
Stan H.
==================